# V. I. Telnov

### v2

@@ 5 Discussion
As one can see, the attainable $L_{e^{+}e^{-}}=(1-2.5)10^{36}$ and $L_{e^{-}e^{-}}=(3-7)10^{36}$ ${\rm\,cm}^{-2}{\rm s}^{-1}$ for reasonable powers of 150 MW and 300 MW for $2502E_{0}=$ 250 and $5002E_{0}=$ 500 GeV, respectively. Cavities at 0.65 GHz allow operation in CW mode at about 2 times lower total power, giving about 2×2$\times$2 × higher luminosity, although they require wider tunnels.
An electron-positron collider is more complicated, as the same bunches must be used many times. This required measures to preserve beam emittance. In fact, it’s like a storage ring with additional components inside, such as linacs and bunch compressors.
Electron-electron ERLC collider is simpler, as each bunch is used only once. Their luminosity is very high, up to $10^{37}$ ${\rm\,cm}^{-2}{\rm s}^{-1}$. Their physics program is also very reach. It’s an attractive option.
@@ 6 Conclusion
Twin $e^{+}e^{-}$ and $e^{-}e^{-}$ linear colliders with the energy recovery open the way to extremely high luminosities. This article demonstrates that the luminosity of the ERLC colliders does not depend on the accelerating gradient and only weakly depends on quality factor of accelerating cavities: $L_{e^{+}e^{-}}\propto Q^{1/2}$, $L_{e^{-}e^{-}}\propto Q^{1/4}$. In the previous paper [5], $e^{+}e^{-}$ ERLC with repetitive use of bunches were considered; this paper considers the case of $e^{-}e^{-}$ ERLC for the first time with single use of electron bunches. The achivable $L_{e^{+}e^{-}}=$ ( 1 – 2.5)×$\times 10^{36}$ for $2E_{0}=$ 250 – 500 GeV and about 3 times higher for $e^{-}e^{-}$ collisions. Such a collider, with energy recovery, can be considered the next (high luminosity) stage of ILC (in its tunnel). Energy recovery superconducting accelerators have many potential applications [11], so we can expect rapid progress in this field.

### v3

@@ 5 Discussion
As one can see, the achievable $L_{e^{+}e^{-}}$ = (1 – 2.5)$10^{36}$ and $L_{e^{-}e^{-}}$ = (3 – 7)$10^{36}$ ${\rm\,cm}^{-2}{\rm s}^{-1}$ for reasonable powers of 150 MW and 300 MW at $2502E_{0}=$ 250 and $5002E_{0}=$ 500 GeV, respectively. Compared to $f_{\rm RF}=1.3$ GHz, cavities operating at 0.65 GHz allow CW operation at about half the power and provide about twice the luminosity, although they require wider tunnels.
An electron-positron collider is more complicated because the same bunches have to be used many times. This requires measures to preserve the emittance of the beam. In fact, it is like a storage ring with additional components inside, such as linear accelerators and bunch compressors.
The ERLC electron-electron collider is simpler because each bunch is used only once (it can be prepared anew each time). Its luminosity can be very high, so it’s also an attractive option.
@@ 6 Conclusion
Twin $e^{+}e^{-}$ and $e^{-}e^{-}$ linear colliders with the energy recovery open the way to extremely high luminosities. This article shows that the luminosity of the ERLC colliders does not depend on the acceleration gradient and only weakly on the quality factor of the RF cavities: $L_{e^{+}e^{-}}\propto Q^{1/2}$, $L_{e^{-}e^{-}}\propto Q^{1/4}$. In the previous paper [5], $e^{+}e^{-}$ ERLC with repeated use of bunches was considered; this paper also considers for the first time the case of $e^{-}e^{-}$ ERLC with single use of electron bunches. The achievable $e^{+}e^{-}$ luminosity $L_{e^{+}e^{-}}=$ ( 1 – 2.5) $10^{36}$ for $2E_{0}=$ 250 – 500 GeV and about 3 times higher in $e^{-}e^{-}$ collisions. Such an energy-recovery collider can be considered as the next high-luminosity stage of the ILC (in its tunnel). Energy-recovery superconducting accelerators have many potential applications [11], so one can expect rapid progress in this field.